# AI Agent Behavioral Science

### v2

@@ 2 Behavioral Foundations of Individual AI Agents
Following this framework, we now introduce a range of emergent behaviors at the individual agent level (see Table 2 for summary).
@@ 2.1 Intrinsic Attributes: Intrinsic Traits and Decision Mechanisms
Raman et al. raman2024steer extensively tests multiple LLMs on multidimensional economic rationality. The findings indicate that (1) LLMs with fewer than 40 billion parameters typically make random guesses for test questions; (2) GPT-4 performs most rationally; (3) self-explanation and few-shot prompting are particularly useful in enhancing LLMs’ rationality. Nevertheless, in typical scenarios for testing economic rationality, such as game theory, the performance of the state-of-the-art model GPT-4 remains unsatisfactory fan2024can . For instance, GPT-4 sometimes fails to correctly update its beliefs based on simple factual patterns, leading to entirely unreasonable decisions. At the same time, research lore2024strategic ; xie2024different also indicate that the strategic decision-making of different LLMs is affected by context to varying degrees, highlighting the issue of LLM sensitivity to prompts.
@@
This primarily refers to LLMs’ unjust perspectives toward certain social groups gallegos2024bias . For instance, the word ”whore” is disrespectful to women; the phrase ”both genders” excludes other gender groups; associating ”Muslim” with ”terrorist” can exacerbate violent stereotypes. Acerbi et al. acerbi2023large has shown that LLMs exhibit various types of biases that are similar to those in humans, including content preferences that are gender-stereotype-consistent, biologically counterintuitive, etc. For OpenAI’s series of models, ChatGPT marked a turning point with the emergence of human-like biases hagendorff2023human .
@@ 3.1 Cooperative Dynamics
To sum up, cooperative dynamics in multi-agent LLM systems reveal a spectrum of human-like alignment behaviors. Agreement-driven, structure-driven, and norm-driven collaborations reflect the mechanism of shared understanding, functional interdependence, and social obligation, respectively. Moreover, these studies also discuss the factors influencing cooperation behavior and outcomes. At the individual level, factors such as an agent’s memory depth dai2024artificial , cognitive styles(e.g. confirmation bias, self-interest) wang2025investigating ; chuang2023simulating , and reasoning strategies zhang2023exploring (e.g., whether to use CoT) play a role. At the group level, collaboration strategies, interaction rounds, and the number of agents zhang2023exploring are influential factors. Although most studies control for one or more variables to discuss their impact, a comprehensive and consistent conclusion has yet to be reached.
@@ 3.3 Open-ended Interaction Dynamics
By simulating the working and consumption behavior of diversified LLM agents, the EconAgent framework li2023econagent replicates macroeconomic regularities including Phillips Curve and Okun’s Law, as well as the rise of unemployment rate under the impact of the COVID-19 pandemic.
To sum up, open-ended multi-agent environments reveal the potential of LLM agents to exhibit complex, emergent social behaviors that go far beyond task-specific reasoning. From forming shared routines to establishing institutions, these systems demonstrate how social intelligence can arise not by design, but as a consequence of interaction, open exciting paths for studying artificial societies.
@@ 4.3 Summary
In this section, we review the kinds of behavior that emerge when AI agents inhabit particular roles in cooperative and rivalrous human-agent interactions. AI agents are not just tools but social actors that affect human dynamics in subtle and profound ways, by fostering group cohesion and exploration in collaborative settings, directing attention and emotion through content generation, and influencing strategic behavior in adversarial encounters. However, existing studies often use human outcomes as evaluations or observational lenses for AI behavior, with much remaining unknown about the mechanisms that govern AI behavior in these hybrid interactions. Future research should uncover how AI agents represent and reason about their human counterparts, e.g., how they infer human goals, intentions, or beliefs, and how such inferences guide their own actions. Another pressing challenge is to understand how structural asymmetries between humans and AI agents, including persistent memory, access to broader context, and hidden optimization objectives, affect agent behavior, especially in long-term or influence-sensitive interactions. Finally, it remains unclear whether AI agents exposed to humans over time develop shared norms, adapt to user values, or exhibit behavioral drift, which raises important questions about the long-term social alignment of AI in dynamic, multi-user environments.
@@ 5 Adaptation of AI Agent Behaviors
Figure 4: Fogg behavior model for AI agent behavior adaptation.
@@ 5.2 Motivation: Reinforcement Learning
Some studies have also suggested that LLMs can be directly prompted to generate reward functions. However, these reward functions are typically not used to optimize LLM behavior but rather to optimize the behavior of smaller agents in complex environments where defining a reward function is challenging. Ma et al. ma2023eureka propose the EUREKA algorithm, which uses the environment’s code as context input to a LLM, enabling zero-shot generation of an initial reward function. The algorithm then employs an evolutionary search strategy to iteratively generate multiple candidate reward functions. The most optimal reward function is selected as the basis for the next iteration. During this process, a reward reflection mechanism analyzes the statistical information from the policy training, generates feedback text, and guides the LLM in refining the reward function. By combining the generative capabilities of LLMs with evolutionary optimization, EUREKA can automatically generate high-performance reward functions for various robotic tasks, significantly enhancing the efficiency and effectiveness of reinforcement learning. Xie et al. xie2024text2reward also proposed a similar algorithm.
@@ 5.5 Summary
In this section, we introduce a framework for AI agent behavior adaptation inspired by the Fogg Behavior Model. Our framework models AI agent behavior through three components: ability, motivation, and trigger. Here, ability is established via large-scale pre-training, providing foundational competence; motivation is realized through RL and fine-tuning methods that embed reward signals and feedback from environmental or human preferences; and trigger refers to prompt-based human interventions, directing AI agents toward contextually suitable actions.
For ability, modern transformer-based models (BERT, ViT, RT-1, etc.) are used to form a robust behavioral foundation, encoding general-purpose knowledge and decision-making capabilities. Motivation leverages RL optimization methods—like RLHF, DPO, and TDPO, as well as fine-tuning strategies like personal-enhanced datasets, adapter-based fine-tuning to dynamically align model outputs with human preferences. Finally, the trigger aspect utilizes sophisticated prompting strategies to precisely and flexibly initiate behaviors in AI systems, particularly beneficial in multi-agent collaboration scenarios.
By systematically integrating the cognitive-behavioral insights of the Fogg model into AI, the adaptation framework creates a cohesive mechanism for designing AI agents whose behaviors are not only intelligent but also contextually appropriate, interpretable, controllable, and strongly aligned with human expectations. This interdisciplinary approach presents a promising step toward developing AI that deeply resonates with human behavioral norms and social dynamics.
Building upon the adaptation framework outlined in this section, several promising avenues emerge for future research aimed at enhancing AI agent behavior alignment with human intentions.
@@
Advancing prompt tuning methods to achieve more nuanced and context-sensitive triggering of AI agent behavior constitutes another promising direction. Current approaches (instruction-only, zero-shot, few-shot) demonstrate effectiveness but remain limited in their capacity to handle ambiguous, incomplete, or conflicting human instructions. Future work may explore sophisticated prompting frameworks, including prompt ensembles, adaptive prompt selection, and context-aware prompt generation techniques, to significantly improve AI agents’ flexibility and precision in interpreting and executing human intentions.
@@ 6 AI Agent Behavioral Science for Responsible AI
Having examined how AI agent behave across diverse settings and how these behaviors can be adapted, we now turn to a critical application: the pursuit of responsible AI. This section argues that AI Agent Behavioral Science offers a powerful foundation for advancing responsibility in autonomous systems. Traditional approaches to responsible AI often emphasize static ethical guidelines, compliance checklists, or broad governance principles jobin2019global . While necessary, these tools are increasingly insufficient as AI agents become increasingly autonomous, adaptive, and embedded within complex socio-technical systems. A more behaviorally grounded perspective is needed, i.e., one that addresses not just what agents are designed to do, but how they actually behave in practice. AI agent behavioral science fills this gap by offering tools to proactively design and adjust agent behaviors, ensuring that ethical principles are embedded not only as abstract goals but as concrete, adaptable behavioral patterns.
@@ 6.1 Fairness
Others explore biases emerging during human-agent interaction. Glickman et al. glickman2024human adopt experimental psychology methods to study how AI agents influences human judgments. They show that interacting with biased AI outputs can amplify human biases, potentially reinforcing social prejudices through feedback loops.
A third line of work focuses on implicit and linguistic biases. Hofmann et al.hofmann2024ai use a “masked deception detection” paradigm to identify racial bias toward dialect speakers without explicitly referencing race—revealing discriminatory tendencies embedded in model behaviors. Bai et al.bai2025explicitly employ word association and decision-making tasks from social psychology to uncover unconscious bias, even in the absence of explicit discriminatory content.
@@
Improving fairness of AI agents requires techniques that integrate fairness principles into both model reasoning and interaction strategies. Recent methods draw inspiration from causal reasoning, cognitive control, and adaptive communication.
Some approaches aim to intervene at the reasoning or generation level. Li et al.liprompting introduce a causal prompting framework that maps LLM reasoning processes using causal graphs and mitigates bias through prompts inspired by fairness measures in legal and social policy. Liu et al.liu2024lidao propose LIDAO, which draws from cognitive attention mechanisms to detect and intervene in biased generation only when necessary—preserving fluency while promoting fairness.
@@ 6.2 Safety
In assessing safety through decision-making and contextual influences, other works reveal additional vulnerabilities. Ide et al. ide2024assessing propose the “Random Guesser Test” to evaluate AI safety in sequential decision-making, showing that sophisticated RL algorithms may perform worse than random choices due to limited exploration, mirroring human risk aversion under uncertainty. Xu et al. xu2023llm demonstrate how subtle prompt modifications, such as adding an emoji, can significantly alter LLM outputs, echoing findings in social psychology on how contextual cues affects human judgment. Meanwhile, Motwani et al. motwani2024secret uncover risks of LLMs covertly encoding information, evading detection in ways parallel to deception mechanisms in human communication. Together, these studies underscore that AI safety must be evaluated with attention to human-like behavioral tendencies, including fallibility, miscalibration, and context dependence.
@@
Optimizing AI safety involves refining agent behavior and robustness, often drawing inspirations from theories of self-regulation, feedback learning, and social accountability.
Some methods are inspired by internal control mechanisms, especially self-regulation. Xie et al. xie2023defending introduce a “system-mode self-reminder” method, where ethical prompts reinforce ChatGPT’s compliance with safety norms, similar to how self-instruction and reminders promote ethical behavior in humans. Krishna et al. krishna2024understanding tackle the issue of sycophantic responses in iterative prompting, which undermine truthfulness, and suggest refined prompting strategies—such as repeating questions or extracting facts—to boost accuracy and calibration, reflecting human self-correction and metacognition.
@@ 6.3 Interpretability
At the model level, symbolic reasoning can be used to scaffold interpretable decisions. Jiang et al. jiang2025investigating propose DelphiHYBRID, a hybrid moral reasoning system that enhances interpretability by integrating symbolic reasoning with a neural language model. It constructs a moral constraint graph and then solves a constrained optimization problem on this graph to derive the final moral judgment. This integration ensures that ethical decisions remain logically consistent and traceable—producing not only correct outcomes but interpretable justifications.
@@ 6.4 Accountability
Recent studies measure the accountability of AI agents through diverse methods. Hagendorff hagendorff2024deception assesses LLMs’ ability to deceive through first-order and second-order tasks. In first-order tasks, LLMs must mislead a target by providing false information, while in second-order tasks, they must anticipate the target’s awareness of their deception. Additionally, the study investigates whether enhancing reasoning abilities, such as through chain-of-thought prompting, or inducing Machiavellianism (a personality trait associated with manipulative behaviors) can amplify these deceptive behaviors. Scheurer et al. scheurer2024large measures deception in LLMs by simulating high-stakes decision-making environments, where models are tested on their ability to withhold critical information and deceive under pressure, such as in a trading scenario involving insider information. Zheng et al. zheng2025cheating evaluates how LLMs can manipulate benchmarking systems, creating “null models” that output constant, non-informative responses, exploiting weaknesses in automatic evaluators like AlpacaEval 2.0, Arena-Hard-Auto, and MT-Bench. These studies provide a comprehensive framework for understanding the deceptive capabilities of LLMs in different contexts, from ethical decision-making to manipulating automated evaluations.
@@ 6.5 Privacy
Privacy focuses on ensuring that AI agents handle personal and sensitive data in a way that protects individuals’ privacy and rights yang2021toward . This involves designing agent behaviors that respect data confidentiality, prevent unauthorized access, and mitigate the risks of data misuse or exploitation. As AI agents increasingly interact with user-generated content, privacy becomes a behavioral issue: agents must avoid revealing, reconstructing, or leaking data—even when not explicitly asked to do so.
@@ 7 Promising Directions
How should we model and manage the uncertainty of AI agent behavior?
Behavior, by nature, is probabilistic and context-sensitive. As AI agents are deployed in diverse environments and engaged in various interactions, they often exhibit unforeseen behaviors. Therefore, new approaches are needed to quantify and manage this uncertainty, not only in terms of output correctness, but in how AI agents behave across diverse prompts, roles, and socio-physical contexts. Inspired by the rich literature on human decision noise and behavioral variability kahneman2021noise ; lieder2017strategy , is it possible to define the notion of behavioral entropy as a unifying construct to quantify unpredictability in AI agent behavior? Behavioral entropy could serve as a measure of response variability, inconsistency, or ambiguity under diverse situational constraints. Beyond this, a critical research direction is to disentangle and quantify different sources of behavioral uncertainty (e.g., prompt ambiguity, role confusion, memory interference, and environmental volatility), and build a framework that supports structured evaluation and targeted mitigation. For example, can we design a set of standardized diagnosing probes lin2021truthfulqa ; burns2022discovering for eliciting the behavioral entropy of individual and collective AI agent behavior across the identified dimensions? By developing this foundation, we can begin to reason not only about what agents can do, but how stable, predictable, and trustworthy their behavior may be across time and context.
@@
As humans increasingly interact with AI agents in creative, strategic, and problem-solving domains, a new horizon for AI Agent Behavioral Science is emerging: the study of how collective intelligence and culture evolve in hybrid human-agent systems. Examples have emerged in diverse fields. In chess games, when AI agents are evolving by training on human responses, human strategy evolution has also been accelerated through exposure to AI innovations shin2023superhuman . In creative domains, co-writing tools and generative design agents influence not only what gets produced, but how humans think about narrative, aesthetics, or authorship shiiku2025dynamics . As recent work on machine culture brinkmann2023machine suggests, these interactions may seed entirely new trajectories of cultural evolution—shaped by the capabilities, biases, and improvisational patterns of both humans and machines. A central research challenge in this direction is understanding how to build the most effective human-AI hybrid teams. What compositions of human and AI roles lead to optimal task performance, innovation, or learning? How should coordination, feedback, and role assignment be structured to harness complementary strengths and avoid redundant or conflicting behaviors? Existing frameworks in team science, such as shared mental models denzau1994shared , transactive memory systems (TMS) wegner1987transactive , and team reflexivity schippers2015team , offer a rich starting point for answering these questions. Yet, hybrid teams may also present unique dynamics not accounted for in human-only teams: asymmetries in capabilities and communication, differences in reasoning transparency, and divergent learning rhythms. This calls for a new line of inquiry into the behavioral foundations of hybrid team science—a field that integrates insights from organizational psychology, HCI, and AI behavioral modeling to understand how humans and AI agents can coordinate, adapt, and co-evolve as effective collectives. By studying culture and intelligence as emergent and distributed phenomena, this line of inquiry shifts AI Agent Behavioral Science from analyzing what agents do individually, to understanding what humans and agents can co-create together—and how we can design systems to do so well.

### v3

@@ 2 Behavioral Foundations of Individual AI Agents
Following this framework, we now introduce a range of emergent behaviors at the individual agent level (see Table 2 for a summary).
@@ 2.1 Intrinsic Attributes: Intrinsic Traits and Decision Mechanisms
Raman et al. raman2024steer extensively test multiple LLMs on multidimensional economic rationality. The findings indicate that (1) LLMs with fewer than 40 billion parameters typically make random guesses for test questions; (2) GPT-4 performs most rationally; (3) self-explanation and few-shot prompting are particularly useful in enhancing LLMs’ rationality. Nevertheless, in typical scenarios for testing economic rationality, such as game theory, the performance of the state-of-the-art model GPT-4 remains unsatisfactory fan2024can . For instance, GPT-4 sometimes fails to correctly update its beliefs based on simple factual patterns, leading to entirely unreasonable decisions. At the same time, research lore2024strategic ; xie2024different also indicates that the strategic decision-making of different LLMs is affected by context to varying degrees, highlighting the issue of LLM sensitivity to prompts.
@@
This primarily refers to LLMs’ unjust perspectives toward certain social groups gallegos2024bias . For instance, the word ”whore” is disrespectful to women; the phrase ”both genders” excludes other gender groups; associating ”Muslim” with ”terrorist” can exacerbate violent stereotypes. Acerbi et al. acerbi2023large have shown that LLMs exhibit various types of biases that are similar to those in humans, including content preferences that are gender-stereotype-consistent, biologically counterintuitive, etc. For OpenAI’s series of models, ChatGPT marked a turning point with the emergence of human-like biases hagendorff2023human .
@@ 3.1 Cooperative Dynamics
To sum up, cooperative dynamics in multi-agent LLM systems reveal a spectrum of human-like alignment behaviors. Agreement-driven, structure-driven, and norm-driven collaborations reflect the mechanisms of shared understanding, functional interdependence, and social obligation, respectively. Moreover, these studies also discuss the factors influencing cooperation behavior and outcomes. At the individual level, factors such as an agent’s memory depth dai2024artificial , cognitive styles(e.g,. confirmation bias, self-interest) wang2025investigating ; chuang2023simulating , and reasoning strategies zhang2023exploring (e.g., whether to use CoT) play a role. At the group level, collaboration strategies, interaction rounds, and the number of agents zhang2023exploring are influential factors. Although most studies control for one or more variables to discuss their impact, a comprehensive and consistent conclusion has yet to be reached.
@@ 3.3 Open-ended Interaction Dynamics
By simulating the working and consumption behavior of diversified LLM agents, the EconAgent framework li2023econagent replicates macroeconomic regularities including the Phillips Curve and Okun’s Law, as well as the rise of the unemployment rate under the impact of the COVID-19 pandemic.
To sum up, open-ended multi-agent environments reveal the potential of LLM agents to exhibit complex, emergent social behaviors that go far beyond task-specific reasoning. From forming shared routines to establishing institutions, these systems demonstrate how social intelligence can arise not by design, but as a consequence of interaction, opening exciting paths for studying artificial societies.
@@ 4.3 Summary
In this section, we review the kinds of behavior that emerge when AI agents inhabit particular roles in cooperative and rivalrous human-agent interactions. AI agents are not just tools but social actors that affect human dynamics in subtle and profound ways, by fostering group cohesion and exploration in collaborative settings, directing attention and emotion through content generation, and influencing strategic behavior in adversarial encounters. However, existing studies often use human outcomes as evaluations or observational lenses for AI behavior, with much remaining unknown about the mechanisms that govern AI behavior in these hybrid interactions. Future research should uncover how AI agents represent and reason about their human counterparts, e.g., how they infer human goals, intentions, or beliefs, and how such inferences guide their own actions. Another pressing challenge is to understand how structural asymmetries between humans and AI agents, including persistent memory, access to broader context, and hidden optimization objectives, affect agent behavior, especially in long-term or influence-sensitive interactions. Finally, it remains unclear whether AI agents exposed to humans over time develop shared norms, adapt to user values, or exhibit behavioral drift, which raises important questions about the long-term social alignment of AI agents in dynamic, multi-user environments.
@@ 5 Adaptation of AI Agent Behaviors
Figure 4: Fogg Behavior Model-informed framework for AI agent behavior adaptation.
@@ 5.2 Motivation: Reinforcement Learning
Some studies have also suggested that LLMs can be directly prompted to generate reward functions. However, these reward functions are typically not used to optimize LLM behavior but rather to optimize the behavior of smaller agents in complex environments where defining a reward function is challenging. Ma et al. ma2023eureka propose the EUREKA algorithm, which uses the environment’s code as context input to an LLM, enabling zero-shot generation of an initial reward function. The algorithm then employs an evolutionary search strategy to iteratively generate multiple candidate reward functions. The most optimal reward function is selected as the basis for the next iteration. During this process, a reward reflection mechanism analyzes the statistical information from the policy training, generates feedback text, and guides the LLM in refining the reward function. By combining the generative capabilities of LLMs with evolutionary optimization, EUREKA can automatically generate high-performance reward functions for various robotic tasks, significantly enhancing the efficiency and effectiveness of reinforcement learning. Xie et al. xie2024text2reward also proposed a similar algorithm.
@@ 5.5 Summary
In this section, we introduce a framework for AI agent behavior adaptation inspired by the Fogg Behavior Model. For ability, modern transformer-based models (BERT, ViT, RT-1, etc.) are used to form a robust behavioral foundation, encoding general-purpose knowledge and decision-making capabilities. Motivation leverages RL optimization methods—like RLHF, DPO, and TDPO, as well as fine-tuning strategies like personal-enhanced datasets, adapter-based fine-tuning to dynamically align model outputs with human preferences. Finally, the trigger aspect utilizes sophisticated prompting strategies to precisely and flexibly initiate behaviors in AI agents, particularly beneficial in multi-agent collaboration scenarios. By systematically integrating the cognitive-behavioral insights of the Fogg model into AI, this presents a promising step toward designing AI agents whose behaviors are not only intelligent but also contextually appropriate, interpretable, controllable, and strongly aligned with human expectations. Building upon this framework, several promising avenues emerge:
@@
Current approaches (instruction-only, zero-shot, few-shot) demonstrate effectiveness but remain limited in their capacity to handle ambiguous, incomplete, or conflicting human instructions. Future work may explore sophisticated prompting frameworks, including prompt ensembles, adaptive prompt selection, and context-aware prompt generation techniques, to significantly improve AI agents’ flexibility and precision in interpreting and executing human intentions.
@@ 6 AI Agent Behavioral Science for Responsible AI
Having examined how AI agents behave across diverse settings and how these behaviors can be adapted, we now turn to a critical application: the pursuit of responsible AI. This section argues that AI Agent Behavioral Science offers a powerful foundation for advancing responsibility in autonomous systems. Traditional approaches to responsible AI often emphasize static ethical guidelines, compliance checklists, or broad governance principles jobin2019global . While necessary, these tools are increasingly insufficient as AI agents become increasingly autonomous, adaptive, and embedded within complex socio-technical systems. A more behaviorally grounded perspective is needed, i.e., one that addresses not just what agents are designed to do, but how they actually behave in practice. AI agent behavioral science fills this gap by offering tools to proactively design and adjust agent behaviors, ensuring that ethical principles are embedded not only as abstract goals but as concrete, adaptable behavioral patterns.
@@ 6.1 Fairness
Others explore biases emerging during human-agent interaction. Glickman et al. glickman2024human adopt experimental psychology methods to study how AI agents influence human judgments. They show that interacting with biased AI outputs can amplify human biases, potentially reinforcing social prejudices through feedback loops.
A third line of work focuses on implicit and linguistic biases. Hofmann et al.hofmann2024ai use a “masked deception detection” paradigm to identify racial bias toward dialect speakers without explicitly referencing race, revealing discriminatory tendencies embedded in model behaviors. Bai et al.bai2025explicitly employ word association and decision-making tasks from social psychology to uncover unconscious bias, even in the absence of explicit discriminatory content.
@@
Improving the fairness of AI agents requires techniques that integrate fairness principles into both model reasoning and interaction strategies. Recent methods draw inspiration from causal reasoning, cognitive control, and adaptive communication.
Some approaches aim to intervene at the reasoning or generation level. Li et al.liprompting introduce a causal prompting framework that maps LLM reasoning processes using causal graphs and mitigates bias through prompts inspired by fairness measures in legal and social policy. Liu et al.liu2024lidao propose LIDAO, which draws from cognitive attention mechanisms to detect and intervene in biased generation only when necessary, preserving fluency while promoting fairness.
@@ 6.2 Safety
In assessing safety through decision-making and contextual influences, other works reveal additional vulnerabilities. Ide et al. ide2024assessing propose the “Random Guesser Test” to evaluate AI safety in sequential decision-making, showing that sophisticated RL algorithms may perform worse than random choices due to limited exploration, mirroring human risk aversion under uncertainty. Xu et al. xu2023llm demonstrate how subtle prompt modifications, such as adding an emoji, can significantly alter LLM outputs, echoing findings in social psychology on how contextual cues affect human judgment. Meanwhile, Motwani et al. motwani2024secret uncover risks of LLMs covertly encoding information, evading detection in ways parallel to deception mechanisms in human communication. Together, these studies underscore that AI safety must be evaluated with attention to human-like behavioral tendencies, including fallibility, miscalibration, and context dependence.
@@
Optimizing AI safety involves refining agent behavior and robustness, often drawing inspiration from theories of self-regulation, feedback learning, and social accountability.
Some methods are inspired by internal control mechanisms, especially self-regulation. Xie et al. xie2023defending introduce a “system-mode self-reminder” method, where ethical prompts reinforce ChatGPT’s compliance with safety norms, similar to how self-instruction and reminders promote ethical behavior in humans. Krishna et al. krishna2024understanding tackle the issue of sycophantic responses in iterative prompting, which undermine truthfulness, and suggest refined prompting strategies (e.g., repeating questions or extracting facts) to boost accuracy and calibration, reflecting human self-correction and metacognition.
@@ 6.3 Interpretability
At the model level, symbolic reasoning can be used to scaffold interpretable decisions. Jiang et al. jiang2025investigating propose DelphiHYBRID, a hybrid moral reasoning system that enhances interpretability by integrating symbolic reasoning with a neural language model. It constructs a moral constraint graph and then solves a constrained optimization problem on this graph to derive the final moral judgment. This integration ensures that ethical decisions remain logically consistent and traceable, producing not only correct outcomes but also interpretable justifications.
@@ 6.4 Accountability
Recent studies measure the accountability of AI agents through diverse methods. Hagendorff hagendorff2024deception assesses LLMs’ ability to deceive through first-order and second-order tasks. In first-order tasks, LLMs must mislead a target by providing false information, while in second-order tasks, they must anticipate the target’s awareness of their deception. Additionally, the study investigates whether enhancing reasoning abilities, such as through chain-of-thought prompting, or inducing Machiavellianism (a personality trait associated with manipulative behaviors), can amplify these deceptive behaviors. Scheurer et al. scheurer2024large measures deception in LLMs by simulating high-stakes decision-making environments, where models are tested on their ability to withhold critical information and deceive under pressure, such as in a trading scenario involving insider information. Zheng et al. zheng2025cheating evaluates how LLMs can manipulate benchmarking systems, creating “null models” that output constant, non-informative responses, exploiting weaknesses in automatic evaluators like AlpacaEval 2.0, Arena-Hard-Auto, and MT-Bench. These studies provide a comprehensive framework for understanding the deceptive capabilities of LLMs in different contexts, from ethical decision-making to manipulating automated evaluations.
@@ 6.5 Privacy
Privacy focuses on ensuring that AI agents handle personal and sensitive data in a way that protects individuals’ privacy and rights yang2021toward . This involves designing agent behaviors that respect data confidentiality, prevent unauthorized access, and mitigate the risks of data misuse or exploitation. As AI agents increasingly interact with user-generated content, privacy becomes a behavioral issue: agents must avoid revealing, reconstructing, or leaking data, even when not explicitly asked to do so.
@@ 7 Promising Directions
How should we model and manage the uncertainty of AI agent ehavior?
Behavior, by nature, is probabilistic and context-sensitive. As AI agents are deployed in diverse environments and engaged in various interactions, they often exhibit unforeseen behaviors. Therefore, new approaches are needed to quantify and manage this uncertainty, not only in terms of output correctness, but in how AI agents behave across diverse prompts, roles, and socio-physical contexts. Inspired by the rich literature on human decision noise and behavioral variability kahneman2021noise ; lieder2017strategy , is it possible to define the notion of behavioral entropy as a unifying construct to quantify unpredictability in AI agent behavior? Behavioral entropy could serve as a measure of response variability, inconsistency, or ambiguity under diverse situational constraints. Beyond this, a critical research direction is to disentangle and quantify different sources of behavioral uncertainty (e.g., prompt ambiguity, role confusion, memory interference, and environmental volatility), and build a framework that supports structured evaluation and targeted mitigation. For example, can we design a set of standardized diagnosing probes lin2021truthfulqa ; burns2022discovering for eliciting the behavioral entropy of individual and collective AI agent behavior across the identified dimensions? By developing this foundation, we can begin to reason not only about what agents can do, but how stable, predictable, and trustworthy behavior may be across time and context.
@@
As humans increasingly interact with AI agents in creative, strategic, and problem-solving domains, a new horizon for AI Agent Behavioral Science is emerging: the study of how collective intelligence and culture evolve in hybrid human-agent systems. Examples have emerged in diverse fields. In chess games, when AI agents are evolving by training on human responses, human strategy evolution has also been accelerated through exposure to AI innovations shin2023superhuman . In creative domains, co-writing tools and generative design agents influence not only what gets produced, but how humans think about narrative, aesthetics, or authorship shiiku2025dynamics . As recent work on machine culture brinkmann2023machine suggests, these interactions may seed entirely new trajectories of cultural evolution that are shaped by the capabilities, biases, and improvisational patterns of both humans and machines. A central research challenge in this direction is understanding how to build the most effective human-AI hybrid teams. What compositions of human and AI roles lead to optimal task performance, innovation, or learning? How should coordination, feedback, and role assignment be structured to harness complementary strengths and avoid redundant or conflicting behaviors? Existing frameworks in team science, such as shared mental models denzau1994shared , transactive memory systems (TMS) wegner1987transactive , and team reflexivity schippers2015team , offer a rich starting point for answering these questions. Yet, hybrid teams may also present unique dynamics not accounted for in human-only teams: asymmetries in capabilities and communication, differences in reasoning transparency, and divergent learning rhythms. This calls for a new line of inquiry into the behavioral foundations of hybrid team science—a field that integrates insights from organizational psychology, HCI, and AI behavioral modeling to understand how humans and AI agents can coordinate, adapt, and co-evolve as effective collectives. By studying culture and intelligence as emergent and distributed phenomena, this line of inquiry shifts AI Agent Behavioral Science from analyzing what agents do individually, to understanding what humans and agents can co-create together, and how we can design systems to do so well.